PHOTOELECTRON SPECTRA FROM WAVE PACKET DYNAMICS

\underline{VADIM MOZHAYSKIY} and ANNA KRYLOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90007.

Franck-Condon factors for photoelectron spectra can be computed from: (1) the overlap between initial and target vibrational wave functions; and (2) Fourier transform of a wave packet time autocorrelation function. These techniques were implemented in the new spectra modeling software using harmonic well approximation and full quantum mechanical treatment. The photoelectron spectrum of N_{3} was modeled using ab-initio potential energy surfaces of the cation electronic states. Anharmonic effects are discussed.